OpenStudy (anonymous):

10x^4-19x^2+6=0 solve

OpenStudy (anonymous):

use radicals as needed and commas

OpenStudy (anonymous):

solve \[10z^2-19z+6=0\] and then replace z by \[x^2\]

OpenStudy (anonymous):

(2x^2- 3 )(5x^2 - 2 ) = 0

OpenStudy (sriram):

substitute x^2=y hence 10y^2-19y+6=0 19+-7 26 or 12 ------= ------- 20 20 x= = or - root of these answers hence we get 4 answers

OpenStudy (anonymous):

\[(2z-3)(5z-2)=0\] \[z=\frac{3}{2}\] or \[z=\frac{2}{5}\]

OpenStudy (anonymous):

x^2 = 1.5 , 0.4 take sq roots to get x

OpenStudy (anonymous):

xo \[x^2=\frac{3}{2}\] \[x=\pm\sqrt{\frac{3}{2}}\] \[x^2=\frac{2}{5}\] \[x=\pm\sqrt{\frac{2}{5}}\]

OpenStudy (anonymous):

\[(\sqrt{2/5},-\sqrt{2/5}),\sqrt{3/2,}-\sqrt{3/2})\]

OpenStudy (anonymous):

would that be correct

OpenStudy (anonymous):

yes